OpenStudy (anonymous):

Convert y=x-4 in standard form.

OpenStudy (saifoo.khan):

standard form = Ax + By = C

OpenStudy (anonymous):

i came up with -x+y= -4. am i incorrect

OpenStudy (mertsj):

You are correct.

OpenStudy (mertsj):

If you did that, Luis, you would get x-y=4

OpenStudy (mertsj):

And I do not believe it is a requirement of general form that A be positive.

OpenStudy (mertsj):